OpenStudy (anonymous):

What is the volume of a cylinder with the following dimensions: Radius 6ft Height 24ft

OpenStudy (apoorvk):

Volume 'V' of cylinder of radius 'r' and height 'h' is \[\large V = \pi r^2h\]

OpenStudy (anonymous):

volume=pi*r^2*h

OpenStudy (lasttccasey):

\[V=πr^2h\]So,\[V=π(6)^2(24)=864π\]

OpenStudy (anonymous):

A. 2712.96 ft3 B. 2712.96 ft2 C. 33.14 ft3 D. 33.14 ft2 these are my choses though, so i'm not really sure?

OpenStudy (anonymous):

at least do the calculation yourself....

OpenStudy (anonymous):

i don't know how to. That's why i'm on this site, Kay, thanks.

OpenStudy (anonymous):

multiply 864pi there's you answer... you can't get more exact than that...

OpenStudy (anonymous):

wait how'd you get that ;o

OpenStudy (apoorvk):

pi = 3.14 or 22/7 ---> multiply this by that 864

OpenStudy (anonymous):

ohhhh okay, yeah, sorry, math is not my subject at all!!!! lol

OpenStudy (lasttccasey):

\[π\approx3.14\]\[864*3.14=2712.96\]So your answer would be,\[A=2712.96ft^3\]Volume is always to the cube root (in this case its ft^3).